# Constraints on Fermionic Dark Matter Absorption from Radiochemical Solar-Neutrino Measurements

### v1

@@ 1 Introduction
In this context, absorption of light fermionic DM provides a well-motivated and experimentally distinctive target. An effective-operator description of fermionic absorption on nuclear targets has been developed in Refs. (Dror et al., 2020b, a), which classifies interactions into neutral-current- and charged-current-like structures and delineates the associated experimental signatures. Because the absorption rate is proportional to the DM number density, absorption searches are naturally sensitive to light DM masses. In the charged-current-like case, precision measurements of solar-neutrino capture rates provide a powerful handle, whereas neutral-current-like signatures are typically constrained via energy-deposition observables in direct-detection experiments (see, e.g., Refs. (Gu and others, 2022; Arnquist and others, 2024; Dai and others, 2022; Adams and others, 2025)).
@@
To enable a direct comparison with existing absorption searches and collider limits on the same effective operators, we adopt the charged-current V–A operator of Refs. (Dror et al., 2020b, a) as our benchmark choice and translate our rate limits onto the standard coupling combination
@@ 2 Analysis Inputs
We use the time-integrated radiochemical production rates measured in the chlorine and gallium programs. For chlorine, we adopt the final Homestake 37Cl result, $R_{\mathrm{Cl}}^{\mathrm{obs}}=2.56\pm 0.16~(\mathrm{stat})\pm 0.16~(\mathrm{syst})~\mathrm{SNU}$ (Cleveland et al., 1998). We combine the statistical and systematic uncertainties in quadrature, yielding $R_{\mathrm{Cl}}^{\mathrm{obs}}=2.56\pm 0.23~\mathrm{SNU}$. For gallium, we use as our default input the weighted average of SAGE, GALLEX, and GNO, $R_{\mathrm{Ga}}^{\mathrm{obs}}=66.1\pm 3.1~\mathrm{SNU}$ (Altmann and others, 2005; Abdurashitov and others, 2009).
@@
and therefore constrain any additional non-negative capture-like contribution once the reaction is kinematically accessible. The kinematic thresholds are $E_{\nu,\mathrm{Cl}}^{\rm th}=0.814~\mathrm{MeV}$ and $E_{\nu,\mathrm{Ga}}^{\rm th}=0.233~\mathrm{MeV}$ (Bahcall, 1997; Bahcall et al., 1998), making gallium sensitive to the low-energy pp-dominated flux in addition to higher-energy components.
The predicted solar-neutrino capture rate depends on the solar-neutrino fluxes $\Phi_{k}$ from the dominant production channels $k=\{\mathrm{pp},\mathrm{pep},{}^{7}\mathrm{Be},{}^{8}\mathrm{B},{}^{13}\mathrm{N},{}^{15}\mathrm{O},{}^{17}\mathrm{F},\mathrm{hep}\}$. The SSM provides self-consistent predictions for these fluxes and their theoretical uncertainties; we adopt the B16 SSM central values and quoted uncertainties (Vinyoles and others, 2017). To make the dominant solar-model systematic explicit, we bracket the metallicity dependence using the high-metallicity B16–GS98 (Grevesse and Sauval, 1998) and low-metallicity B16–AGSS09met (Asplund et al., 2009) realizations.
For each target $T\in\{\mathrm{Cl},\mathrm{Ga}\}$, the capture-rate prediction depends on the charged-current capture cross section $\sigma_{T}(E)$. We adopt $\sigma_{T}(E)$ for 37Cl and 71Ga (with quoted uncertainties) from the Bahcall et al. calculations and subsequent updates (Bahcall et al., 1996; Bahcall, 1997; Bahcall et al., 1998), and evaluate the individual rate contributions by numerical integration over the standard solar spectra.
@@
where $\phi_{k}(E)$ is the normalized energy spectrum of component $k$ (including line spectra for monochromatic sources). We evaluate $P_{ee}(E)$ in a three-flavor Mikheyev–Smirnov–Wolfenstein (MSW) framework using an adiabatic approximation (Wolfenstein, 1978; Mikheyev and Smirnov, 1985; Parke, 1986; Kuo and Pantaleone, 1989).
We incorporate external constraints on the oscillation parameters entering $P_{ee}(E)$. To avoid potential circularity with solar-neutrino rate information, we take the “solar” parameters $(\sin^{2}\theta_{12},\Delta m_{21}^{2})$ from reactor $\bar{\nu}_{e}$ oscillation measurements. Concretely, we adopt the first JUNO reactor result (Normal Ordering) (Abusleme and others, 2025): $\sin^{2}\theta_{12}=0.3092\pm 0.0087$ and $\Delta m_{21}^{2}=(7.50\pm 0.12)\times 10^{-5}\ \mathrm{eV}^{2}$. For $\theta_{13}$ we use the Particle Data Group (PDG) world average (Navas and others, 2024): $\sin^{2}\theta_{13}=(2.19\pm 0.07)\times 10^{-2}$. For solar $\nu_{e}$ survival probabilities, $\theta_{23}$ and $\delta_{\rm CP}$ do not enter at leading order and are neglected.
@@ 3 Analysis Framework
We infer upper bounds on additional, non-negative capture-like contributions by comparing radiochemical production-rate measurements to the solar $\nu_{e}$ capture. The capture-rate prediction depends on nuisance parameters describing solar-neutrino fluxes, neutrino oscillations, and capture cross-section inputs. We treat the DM-induced contribution either (i) phenomenologically as independent additive rates (in SNU) for each target, or (ii) in our charged-current absorption interpretation, parameterized by $y\equiv m_{\chi}^{2}/(4\pi\Lambda^{4})$ (Dror et al., 2020b, a). Throughout, we adopt flat priors on the non-negative DM signal parameters and compute one-sided Bayesian upper limits at 90% credible level.
@@
where $p_{e}=\sqrt{E_{e}^{2}-m_{e}^{2}}$ and $E_{e}=m_{e}+(m_{\chi}-E_{\nu,T}^{\rm th})$. The factor $1/2$ in Eq. (11) follows the convention of Refs. (Dror et al., 2020b, a).
@@
$p(R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}},R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}}\,|\,\mathrm{data})\propto\mathcal{L}_{\mathrm{marg}}(R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}},R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}})\,\pi(R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}})\,\pi(R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}})\,.$ (19)
We obtain one-sided 90% credible upper limits from the 1D marginalized posteriors, e.g.
$\int_{0}^{R_{\chi,T,90}^{\mathrm{(free)}}}dR_{\chi,T}^{\mathrm{(free)}}\;p(R_{\chi,T}^{\mathrm{(free)}}\,|\,\mathrm{data})=0.9\,,\qquad T\in\{\mathrm{Cl},\mathrm{Ga}\}\,.$ (20)
@@ 4 Results
In this section we present constraints on additional non-negative capture-like contributions from fermionic DM absorption using radiochemical solar-neutrino measurements. We first report rate-level (model-independent) upper limits on additive contributions in SNU, and then present our main charged-current limits, in terms of $y=m_{\chi}^{2}/(4\pi\Lambda^{4})=\kappa m_{\chi}^{2}/(4\pi)$ (Dror et al., 2020b, a). Results are shown for both B16 SSM realizations (GS98 and AGSS09met), which bracket the dominant metallicity-driven flux systematic.
@@ 4.2 Main limits in charged-current absorption
Figure 2 presents our primary constraints on charged-current absorption, shown as a 90% credible upper bound on $y=m_{\chi}^{2}/(4\pi\Lambda^{4})$ as a function of the DM mass. The curves start at $m_{\chi}\gtrsim E_{\nu,\mathrm{Ga}}^{\rm th}\simeq 0.233~\mathrm{MeV}$, reflecting the kinematic threshold for charged-current capture on 71Ga; below threshold the absorption-induced transition is forbidden. Once $m_{\chi}\gtrsim E_{\nu,\mathrm{Cl}}^{\rm th}\simeq 0.814~\mathrm{MeV}$, the chlorine channel opens and rapidly dominates the combined likelihood.
Figure 2 also overlays representative external constraints on the charged-current absorption, including the EXO-200 charged-current absorption search (Al Kharusi and others, 2023) and the indirect/collider limits compiled in Refs. (Dror et al., 2020b, a). The collider/decay curves should be regarded as benchmark interpretations: they rely on the validity of an effective field theory (EFT) description at the relevant momentum transfers, or equivalently on assumptions about the ultraviolet (UV) completion (mediator mass/width and couplings) that generates the low-energy operator. The radiochemical limits are complementary: they reach down to the gallium threshold ($m_{\chi}\simeq 0.233~\mathrm{MeV}$) and, above the chlorine threshold, become driven by the precisely measured 37Cl rate.
Figure 2: Main 90% credible upper limits on charged-current fermionic DM absorption, shown as a bound on $y=m_{\chi}^{2}/(4\pi\Lambda^{4})$ as a function of the DM mass. Curves correspond to our baseline “pep-normalized” operator mapping, in which the effective nuclear response is calibrated to tabulated pep $\nu_{e}$ capture cross sections for each target. Results are shown for the two B16 SSM realizations (B16–GS98 and B16–AGSS09met), which bracket the dominant metallicity-driven flux systematic. For context we overlay representative external constraints, including the EXO-200 absorption limit (Al Kharusi and others, 2023) and the collider/decay bounds as compiled in Refs. (Dror et al., 2020b, a).
@@ 5 Discussion
Interpreting a rate constraint in terms of $y$ requires an effective absorption kernel $A_{T}(m_{\chi})$ (Eq. 12) and introduces an operator-mapping uncertainty associated with nuclear response modeling. We minimize this additional uncertainty by adopting a pep-normalized operator mapping, in which the dominant target-dependent nuclear-structure dependence is absorbed into a single effective matrix element $M_{\rm eff,T}^{2}$ calibrated to tabulated pep $\nu_{e}$ capture inputs. Residual dependence on Coulomb corrections, recoil effects, and final-state energy dependence is expected to be subdominant at the level of current uncertainties. This operator-mapping uncertainty is conceptually distinct from the solar-metallicity systematic in the SSMs, which we represent explicitly with the B16–GS98 and B16–AGSS09met realizations (Vinyoles and others, 2017; Grevesse and Sauval, 1998; Asplund et al., 2009).
The resulting radiochemical limits are complementary to other probes of fermionic DM absorption. Xenon-based searches such as EXO-200 constrain the same charged-current absorption parameterization ($y$) at higher masses (Al Kharusi and others, 2023), while collider and decay bounds can be interpreted in the same benchmark framework (Dror et al., 2020b, a). Recently, Ref. (Richardson and others, 2025) studied the prospective sensitivity of nEXO to the charged-current absorption channel, exploiting a delayed-coincidence signature from low-lying $^{136}\mathrm{Cs}$ isomers.
@@ 6 Conclusion
We derive Bayesian upper limits on fermionic DM absorption using time-integrated radiochemical solar-$\nu_{e}$ capture-rate measurements on 37Cl and 71Ga. Our likelihood combines the Homestake chlorine rate (Cleveland et al., 1998) with the weighted gallium average of SAGE, GALLEX, and GNO (Altmann and others, 2005; Abdurashitov and others, 2009), and marginalizes over the dominant theory uncertainties from B16 solar fluxes (Vinyoles and others, 2017) (bracketed by GS98 and AGSS09met (Grevesse and Sauval, 1998; Asplund et al., 2009)), capture cross sections $\sigma_{T}(E)$ and their uncertainties (Bahcall et al., 1996; Bahcall, 1997; Bahcall et al., 1998), and externally constrained oscillation parameters (Abusleme and others, 2025; Navas and others, 2024).
As a rate-level (model-independent) intermediate result, we constrain non-negative additive capture-like rates $R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}}$ and $R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}}$ (SNU) with the combined chlorine-and-gallium likelihood; above the chlorine threshold the $R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}}$ limit is driven by the small fractional uncertainty of Homestake. We then interpret the constraints in the charged-current absorption as 90% credible upper bounds on $y=m_{\chi}^{2}/(4\pi\Lambda^{4})=\kappa m_{\chi}^{2}/(4\pi)$ from the gallium threshold at $m_{\chi}\simeq 0.233~\mathrm{MeV}$ up to tens of MeV. Our limits are particularly robust in this benchmark because the operator mapping is pep-normalized, i.e., anchored to solar $\nu_{e}$ capture inputs on the same targets. These radiochemical limits provide a simple, data-driven complement to xenon-based absorption searches such as EXO-200 (Al Kharusi and others, 2023) and to collider/decay interpretations in the same benchmark (subject to EFT/UV assumptions) (Dror et al., 2020b, a), and show that existing solar radiochemical data probe a broad and well-motivated portion of charged-current absorption parameter space.

### v2

@@ 1 Introduction
In this context, absorption of light fermionic DM provides a well-motivated and experimentally distinctive target. An effective-operator description of fermionic absorption on nuclear targets has been developed in Refs. (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ), which classifies interactions into neutral-current- and charged-current-like structures and delineates the associated experimental signatures. Because the absorption rate is proportional to the DM number density, absorption searches are naturally sensitive to light DM masses. In the charged-current-like case, precision measurements of solar-neutrino capture rates provide a powerful handle. More broadly, absorption leads to a characteristic energy deposition, so complementary constraints can be obtained from direct-detection searches for absorption signals on nuclear and electron targets (Gu:2022PandaX4T, ; Arnquist:2024MajoranaDM, ; Dai:2022CDEX10, ; Adams:2025PICO60, ; PandaX:2022ood, ; Ge:2022ius, ).
@@
It is worth noting that the observed chlorine and gallium radiochemical production rates are broadly consistent with the corresponding solar-neutrino capture-rate expectations once the electron-neutrino survival probability is evaluated using oscillation parameters constrained independently by reactor antineutrino experiments. This agreement implies that only limited room remains for additional non-negative capture-like contributions, which naturally motivates reinterpreting the radiochemical datasets as sensitive probes of fermionic DM absorption.
@@
To enable a direct comparison with existing absorption searches and collider limits on the same effective operators, we adopt the charged-current V–A operator of Refs. (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ) as our benchmark choice and translate our rate limits onto the standard coupling combination
@@ 2 Analysis Inputs
We use the time-integrated radiochemical production rates measured in the chlorine and gallium programs. For chlorine, we adopt the final Homestake 37Cl result, $R_{\mathrm{Cl}}^{\mathrm{obs}}=2.56\pm 0.16~(\mathrm{stat})\pm 0.16~(\mathrm{syst})~\mathrm{SNU}$ (Cleveland:1998, ). We combine the statistical and systematic uncertainties in quadrature, yielding $R_{\mathrm{Cl}}^{\mathrm{obs}}=2.56\pm 0.23~\mathrm{SNU}$. For gallium, we use as our default input the weighted average of SAGE, GALLEX, and GNO, $R_{\mathrm{Ga}}^{\mathrm{obs}}=66.1\pm 3.1~\mathrm{SNU}$ (Altmann:2005, ; Abdurashitov:2009, ).
@@
and therefore constrain any additional non-negative capture-like contribution once the reaction is kinematically accessible. The kinematic thresholds are $E_{\nu,\mathrm{Cl}}^{\rm th}=0.814~\mathrm{MeV}$ and $E_{\nu,\mathrm{Ga}}^{\rm th}=0.233~\mathrm{MeV}$ (Bahcall:1997GaXS, ; Bahcall:1998Uncertainties, ), making gallium sensitive to the low-energy pp-dominated flux in addition to higher-energy components.
The predicted solar-neutrino capture rate depends on the solar-neutrino fluxes $\Phi_{k}$ from the dominant production channels $k=\{\mathrm{pp},\mathrm{pep},{}^{7}\mathrm{Be},{}^{8}\mathrm{B},{}^{13}\mathrm{N},{}^{15}\mathrm{O},{}^{17}\mathrm{F},\mathrm{hep}\}$. The SSM provides self-consistent predictions for these fluxes and their theoretical uncertainties; we adopt the B16 SSM central values and quoted uncertainties (B16SSM, ). To make the dominant solar-model systematic explicit, we bracket the metallicity dependence using the B16–GS98 (GS98, ) and B16–AGSS09met (AGSS09, ) realizations. Here, B16-GS98 and B16-AGSS09met correspond to the high- and low-metallicity realizations of the SSM, respectively. The difference arises from the heavy-element abundances adopted in the solar composition, which affect the solar interior opacity and therefore the predicted solar-neutrino fluxes. In practice, the higher-metallicity B16-GS98 model generally predicts slightly larger solar-neutrino capture rates for the radiochemical targets considered here. Consequently, less room remains for an additional non-negative DM-induced contribution, which leads to stronger limits than those obtained with the lower-metallicity B16-AGSS09met model.
For each target $T\in\{\mathrm{Cl},\mathrm{Ga}\}$, the capture-rate prediction depends on the charged-current capture cross section $\sigma_{T}(E)$. We adopt $\sigma_{T}(E)$ for 37Cl and 71Ga (with quoted uncertainties) from the Bahcall et al. calculations and subsequent updates (Bahcall:1996B8Spec, ; Bahcall:1997GaXS, ; Bahcall:1998Uncertainties, ), and evaluate the individual rate contributions by numerical integration over the standard solar spectra.
@@
where $\phi_{k}(E)$ is the normalized energy spectrum of component $k$ (including line spectra for monochromatic sources). We evaluate $P_{ee}(E)$ in a three-flavor Mikheyev–Smirnov–Wolfenstein (MSW) framework using an adiabatic approximation (Wolfenstein:1978, ; MSW:1985, ; Parke:1986, ; KuoPantaleone:1989, ).
We incorporate external constraints on the oscillation parameters entering $P_{ee}(E)$. To avoid potential circularity with solar-neutrino rate information, we take the “solar” parameters $(\sin^{2}\theta_{12},\Delta m_{21}^{2})$ from reactor $\bar{\nu}_{e}$ oscillation measurements. Concretely, we adopt the first JUNO reactor result (Normal Ordering) (JUNOFirstOsc, ): $\sin^{2}\theta_{12}=0.3092\pm 0.0087$ and $\Delta m_{21}^{2}=(7.50\pm 0.12)\times 10^{-5}\ \mathrm{eV}^{2}$. For $\theta_{13}$ we use the Particle Data Group (PDG) world average (PDG2024, ): $\sin^{2}\theta_{13}=(2.19\pm 0.07)\times 10^{-2}$. For solar $\nu_{e}$ survival probabilities, $\theta_{23}$ and $\delta_{\rm CP}$ do not enter at leading order and are neglected.
@@ 3 Analysis Framework
We infer upper bounds on additional, non-negative capture-like contributions by comparing radiochemical production-rate measurements to the solar $\nu_{e}$ capture. The capture-rate prediction depends on nuisance parameters describing solar-neutrino fluxes, neutrino oscillations, and capture cross-section inputs. We treat the DM-induced contribution either (i) phenomenologically as independent additive rates (in SNU) for each target, or (ii) in our charged-current absorption interpretation, parameterized by $y\equiv m_{\chi}^{2}/(4\pi\Lambda^{4})$ (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ). Throughout, we adopt flat priors on the non-negative DM signal parameters and compute one-sided Bayesian upper limits at 90% credible level.
@@
The superscript “(free)” indicates that these parameters are treated as model-independent free variables representing non-negative additive capture-like rates expressed in SNU, prior to imposing the charged-current absorption mapping used in the benchmark operator interpretation discussed below.
@@
where $p_{e}=\sqrt{E_{e}^{2}-m_{e}^{2}}$ and $E_{e}=m_{e}+(m_{\chi}-E_{\nu,T}^{\rm th})$. The factor $1/2$ in Eq. (11) follows the convention of Refs. (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ).
@@
$p(R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}},R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}}\,|\,\mathrm{data})\propto\mathcal{L}_{\mathrm{marg}}(R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}},R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}})\,\pi(R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}})\,\pi(R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}})\,$, (19)
where $\pi$ denotes the prior probability density. In particular, $\pi(\boldsymbol{\eta})$ represents the joint prior for the nuisance parameters included in the analysis, while $\pi(R_{\chi,T}^{\mathrm{(free)}})$ denotes the prior for the dark-matter-induced capture-like rate parameters with $T\in\{\mathrm{Cl},\mathrm{Ga}\}$. We obtain one-sided 90% credible upper limits from the 1D marginalized posteriors, e.g.
$\int_{0}^{R_{\chi,T,90}^{\mathrm{(free)}}}dR_{\chi,T}^{\mathrm{(free)}}\;p(R_{\chi,T}^{\mathrm{(free)}}\,|\,\mathrm{data})=0.9\,$, (20)
@@ 4 Results
In this section we present constraints on additional non-negative capture-like contributions from fermionic DM absorption using radiochemical solar-neutrino measurements. We first report rate-level (model-independent) upper limits on additive contributions in SNU, and then present our main charged-current limits, in terms of $y=m_{\chi}^{2}/(4\pi\Lambda^{4})=\kappa m_{\chi}^{2}/(4\pi)$ (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ). Results are shown for both B16 SSM realizations (GS98 and AGSS09met), which bracket the dominant metallicity-driven flux systematic.
@@ 4.2 Main limits in charged-current absorption
Figure 2 presents our primary constraints on charged-current absorption, shown as a 90% credible upper bound on $y=m_{\chi}^{2}/(4\pi\Lambda^{4})$ as a function of the DM mass. The curves start at $m_{\chi}\gtrsim E_{\nu,\mathrm{Ga}}^{\rm th}\simeq 0.233~\mathrm{MeV}$, reflecting the kinematic threshold for charged-current capture on 71Ga; below threshold the absorption-induced transition is forbidden. Once $m_{\chi}\gtrsim E_{\nu,\mathrm{Cl}}^{\rm th}\simeq 0.814~\mathrm{MeV}$, the chlorine channel opens and rapidly dominates the combined likelihood. The difference between the limits obtained using B16-GS98 and B16-AGSS09met directly reflects the metallicity dependence of the underlying solar-neutrino flux predictions: larger predicted capture rates leave less room for additional DM-induced contributions and therefore yield stronger upper limits.
Figure 2 also overlays representative external constraints on the charged-current absorption, including the EXO-200 charged-current absorption search (EXO-200:2022adi, ) and the indirect/collider limits compiled in Refs. (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ). The collider/decay curves should be regarded as benchmark interpretations: they rely on the validity of an effective field theory (EFT) description at the relevant momentum transfers, or equivalently on assumptions about the ultraviolet (UV) completion (mediator mass/width and couplings) that generates the low-energy operator. The radiochemical limits are complementary: they reach down to the gallium threshold ($m_{\chi}\simeq 0.233~\mathrm{MeV}$) and, above the chlorine threshold, become driven by the precisely measured 37Cl rate.
Figure 2: Main 90% credible upper limits on charged-current fermionic DM absorption, shown as a bound on $y=m_{\chi}^{2}/(4\pi\Lambda^{4})$ as a function of the DM mass. Curves correspond to our baseline “pep-normalized” operator mapping, in which the effective nuclear response is calibrated to tabulated pep $\nu_{e}$ capture cross sections for each target. Results are shown for the two B16 SSM realizations (B16–GS98 and B16–AGSS09met), which bracket the dominant metallicity-driven flux systematic. For context we overlay representative external constraints, including the EXO-200 absorption limit (EXO-200:2022adi, ) and the collider/decay bounds as compiled in Refs. (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ).
@@ 5 Discussion
Interpreting a rate constraint in terms of $y$ requires an effective absorption kernel $A_{T}(m_{\chi})$ (Eq. 12) and introduces an operator-mapping uncertainty associated with nuclear response modeling. We minimize this additional uncertainty by adopting a pep-normalized operator mapping, in which the dominant target-dependent nuclear-structure dependence is absorbed into a single effective matrix element $M_{\rm eff,T}^{2}$ calibrated to tabulated pep $\nu_{e}$ capture inputs. Residual dependence on Coulomb corrections, recoil effects, and final-state energy dependence is expected to be subdominant at the level of current uncertainties. This operator-mapping uncertainty is conceptually distinct from the solar-metallicity systematic in the SSMs, which we represent explicitly with the B16–GS98 and B16–AGSS09met realizations (B16SSM, ; GS98, ; AGSS09, ).
The resulting radiochemical limits are complementary to other probes of fermionic DM absorption. Xenon-based searches such as EXO-200 constrain the same charged-current absorption parameterization ($y$) at higher masses (EXO-200:2022adi, ), while collider and decay bounds can be interpreted in the same benchmark framework (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ). Recently, Ref. (Richardson:2025nEXO, ) studied the prospective sensitivity of nEXO to the charged-current absorption channel, exploiting a delayed-coincidence signature from low-lying $^{136}\mathrm{Cs}$ isomers.
@@ 6 Conclusion
We derive Bayesian upper limits on fermionic DM absorption using time-integrated radiochemical solar-$\nu_{e}$ capture-rate measurements on 37Cl and 71Ga. Our likelihood combines the Homestake chlorine rate (Cleveland:1998, ) with the weighted gallium average of SAGE, GALLEX, and GNO (Altmann:2005, ; Abdurashitov:2009, ), and marginalizes over the dominant theory uncertainties from B16 solar fluxes (B16SSM, ) (bracketed by GS98 and AGSS09met (GS98, ; AGSS09, )), capture cross sections $\sigma_{T}(E)$ and their uncertainties (Bahcall:1996B8Spec, ; Bahcall:1997GaXS, ; Bahcall:1998Uncertainties, ), and externally constrained oscillation parameters (JUNOFirstOsc, ; PDG2024, ).
As a rate-level (model-independent) intermediate result, we constrain non-negative additive capture-like rates $R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}}$ and $R_{\chi,\mathrm{Ga}}^{\mathrm{(free)}}$ (SNU) with the combined chlorine-and-gallium likelihood; above the chlorine threshold the $R_{\chi,\mathrm{Cl}}^{\mathrm{(free)}}$ limit is driven by the small fractional uncertainty of Homestake. We then interpret the constraints in the charged-current absorption as 90% credible upper bounds on $y=m_{\chi}^{2}/(4\pi\Lambda^{4})=\kappa m_{\chi}^{2}/(4\pi)$ from the gallium threshold at $m_{\chi}\simeq 0.233~\mathrm{MeV}$ up to tens of MeV. Our limits are particularly robust in this benchmark because the operator mapping is pep-normalized, i.e., anchored to solar $\nu_{e}$ capture inputs on the same targets. These radiochemical limits provide a simple, data-driven complement to xenon-based absorption searches such as EXO-200 (EXO-200:2022adi, ) and to collider/decay interpretations in the same benchmark (subject to EFT/UV assumptions) (Dror:2020FDM_PRL, ; Dror:2020FDM_JHEP, ), and show that existing solar radiochemical data probe a broad and well-motivated portion of charged-current absorption parameter space.